the percentage composition of a gas is :
Nitrogen 82.35%,Hydrogen 17.64%
Find the empirical formula of the gas.
 

Grade:10

1 Answers

Ravleen Kaur
6 years ago
Hello Student,
Given:
Mass percentage of nitrogen is N% = 82.35
Mass percentage of Hydrogen is H% = 17.64

let the mass of the compound be 100g
Then no of moles of nitrogen is = N%/14 = 82.35/14 =5.88214
no of moles of hydrgen is = H%/1 = 17.64

Approximation of hydrogen and nitrogen moles to nearest integer (17.64/5.88214 =3)
we obtain
the empirical formula NH3

Regards,
IHE